OpenStudy (anonymous):

approximately how many square miles are there in 25 square kilometers ? round to the nearest tenth

OpenStudy (jdoe0001):

how many miles in 1 kilometer anyway? or how many kms in 1 mile?

OpenStudy (anonymous):

0.621371

OpenStudy (jdoe0001):

yeap so there are 0.621371 miles in 1 km so

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what gd are u in anyways

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} 0.621371m&1km\\ (0.621371m)^2&(1km)^2\\ 0.621372m^2&1km^2\\ \hline\\ 25\cdot 0.621372m^2& 25km^2 \end{array}\)

OpenStudy (jdoe0001):

woops missed something there one sec

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} 0.621371m&1km\\ (0.621371m)^2&(1km)^2\\ 0.621372^2m^2&1km^2\\ \hline\\ 25\cdot 0.621372^2m^2& 25km^2 \end{array}\)

OpenStudy (jdoe0001):

hmmm I ended up with 372 for whatever reason lol guess is just a typo

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} 0.621371m&1km\\ (0.621371m)^2&(1km)^2\\ 0.621371^2m^2&1km^2\\ \hline\\ 25\cdot 0.621371^2m^2& 25km^2 \end{array}\) there

OpenStudy (anonymous):

lol so you r in 11th gd jw

OpenStudy (jdoe0001):

\(\huge ☺ \)

OpenStudy (anonymous):

lol so you are in the 11th grade

OpenStudy (anonymous):

lol i just wanted to know what school u go too